# dMel: Speech Tokenization made Simple

### v2

@@ Abstract
Large language models have revolutionized natural language processing by leveraging self-supervised pretraining on vast textual data. Inspired by this success, researchers have investigated complicated speech tokenization methods to discretize continuous speech signals so that language modeling techniques can be applied to speech data. However, existing approaches either model semantic (content) tokens, potentially losing acoustic information, or model acoustic tokens, risking the loss of semantic (content) information. Having multiple token types also complicates the architecture and requires additional pretraining. Here we show that discretizing mel-filterbank channels into discrete intensity bins produces a simple representation (dMel), that performs better than other existing speech tokenization methods. Using an LM-style transformer architecture for speech-text modeling, we comprehensively evaluate different speech tokenization methods on speech recognition (ASR) and speech synthesis (TTS). Our results demonstrate the effectiveness of dMel in achieving high performance on both tasks within a unified framework, paving the way for efficient and effective joint modeling of speech and text. The code will be open sourced.
@@ 1 Introduction
Current speech tokenization approaches can be broadly categorized into two types: semantic (content) tokens and acoustic tokens [^1] We use a word ‘semantic’ with the meaning of ‘content’ to keep prior work notation [7].. Semantic tokens, extracted from self-supervised (SSL) pretrained speech models [2, 16], where the speech signal is first encoded into speech representations and then clustered into semantic tokens with $k$-means method. However, such SSL pretrained models are not useful for high fidelity speech synthesis as speaker identity and other details of raw speech are lost in training [7]. Conversely, acoustic tokens can be obtained from audio compression models that are trained to compress the speech signal into codebook indices with residual vector quantization (RVQ) and reconstruction objectives [42, 11]. These tokens prioritize acoustic reconstruction but lose semantic information which can lead to poorer results in generating audio [38].
@@
We find that the word error rate (WER) of an ASR system run on the reconstructed waveforms, is quite similar to the WER of the same system run on the ground-truth audio, showing that dMel captures the acoustic information needed to reconstruct good waveforms. Similarly, we find that the WER of ASR models trained on Mel and dMel are similar, indicating that dMel are good at preserving semantic content that can easily be found by ASR models. This shows that discretizing Mel has limited impact on information content.
@@
By operating on the log mel-filterbanks and preserving the frequency and intensity information (with some loss of resolution from discretization), dMel inherently preserves both semantic and acoustic information in a unified representation, without the need for separate tokenization or additional pretraining of a tokenization model. There are many advantages to discretizing log mel-filterbanks:
  • Log mel-filterbanks is an interpretable representation of speech signal, where both the semantic and acoustic information is preserved. As discretization has little impact, dMel inherits the properties.
  • dMel is a model-free representation grounded in raw acoustic space. As a result it can be converted to waveforms by any mel-filterbank vocoder, unlike other tokenization schemes that have feature representations that are intricately coupled to both the encoder and the decoder.
  • Different channels of dMel do not have the complex hierarchical dependencies on each other that is typical of coarse-to-fine acoustic tokens; we find that they can be modeled independently in each frame using a simple decoder-only (LM-style) transformer architecture.
@@
Through comprehensive evaluations, we show that using dMel allows us to use a single decoder-only model, and achieve high performance on both automatic speech recognition (ASR) and text-to-speech (TTS) tasks. The ASR task validates that dMel preserves semantic information, while the TTS task shows that dMel are useful for high-fidelity acoustic reconstruction of speech. We also compare dMel to other tokenization methods and find that dMel achieves the best WER for ASR task, which indicates that the semantic information is well preserved. Also, dMel achieves the lower WER score for TTS task when using WhisperX [4] for automatic evaluation and we find model trained with dMel can generate long and natural speech samples (see Supplementary Material).
@@
Figure 1: Prior works on speech tokenization use either heavy self-supervised pretrained encoders [2, 16] to extract semantic tokens (and train a separate decoder for it [22]) or learn compression encoder-decoder models with residual vector quantizations [42, 11] to obtain acoustic tokens. By contrast we eliminate the encoder and simply discretize mel-filerbanks (dMel) to encode audio, and use a simple mel-filterbank vocoder [41] to reconstruct speech signals.
@@ 2 Method
2.2 Unified Speech-Text Transformer Decoder
Figure 2: Unified Speech-Text Transformer Decoder with speech tokens as dMel.
@@ 2.2 Unified Speech-Text Transformer Decoder
Table 2: Comparison between different speech tokenizers: dMel (ours), HuBERT-LM and SpeechTokenizer. For dMel we use $N=80$ log mel-filterbanks and $162^{K}=$ 16 values of the codebook $\mathbf{C}$.
dMel HuBERT-KM SpeechTokenizer
Codebook Size 16 200 1024
Code Dimension 80 1 8
Training-free?
Speaker Representation
To properly model multi-speaker data, we also include speaker embeddings as input to the transformer decoder. The speaker embeddings are extracted from an independent dvector [37] model [^3] We use a pretrained model “Speaker Encoder” from the YourTTS [9] repository https://github.com/Edresson/YourTTS.. We use a learnable linear layer to map the speaker embeddings to the same dimension as the speech and text token embeddings $D$. The speaker representation is optional for ASR task, but required for TTS task. Hence, during the training, it is applied for text-to-speech and ignored for speech-to-text.
@@
The transformer decoder is trained end-to-end on a combined dataset of speech and text pairs. For TTS training, the input sequence is constructed by concatenating the speaker embedding (extracted from a random audio for the same speaker of the current sample), text tokens, and speech tokens. For ASR training, the input sequence is constructed by concatenating the speech tokens and text tokens. Both tasks are trained with causal masking, where the model is trained to predict the next token based on the previous tokens. The loss is calculated using the cross-entropy loss between the predicted tokens and the ground-truth tokens. Loss calculation is skipped on the speech tokens for ASR task and on the text tokens for TTS task. Note, that all frequency channels at time frame $t$ for dMel tokenizer are predicted independently and in parallel.
To capture the relative distances between tokens in the input sequence, we apply multiplicative relative positional embedding RoPE [35]. This allows the model to learn the positional relationships between speech tokens, text tokens, and speaker embeddings, enhancing its ability to generate coherent output sequences. For positional embeddings we do not distinguish between text, speech and speaker tokens and thus having global positions notation across all of them, see Figure 2.
Robust Training
@@ 3 Experiments
3.1 Training Data
We use several open-sourced datasets with paired speech and text transcription to conduct experiments: i) LibriSpeech [25] dataset (CC BY 4.0) consists of English speech recordings (960h, 16kHz) from various speakers ($\sim$2k) and conditions; ii) LibriTTS [43] (CC BY 4.0) dataset (500h) derived from LibriSpeech improves on it with the proper sentence split, text normalization and keeping samples 24kHz; iii) VCTK [40] contains 44h of English speech (108 speakers); iv) LJSpeech [17] (public domain in US) is a single speaker English audio recordings of 16kHz with read speech from LibriVox [^4] https://librivox.org/pages/public-domain/.. While LibriSpeech is used to train ASR and TTS models, LibriTTS, VCTK and LJSpeech are only used to train the TTS.
3.2 Training Configuration
3.3 Main Results
3.3.1 Speech Reconstruction
Following [46], we randomly sample 300 speech utterances and their ground truth transcriptions from the LibriSpeech test-clean dataset. We use the speech2unit and unit2speech modules to convert the speech signal to speech tokens and then reconstruct the speech signal from the speech tokens. We compute the WER between the ASR outputs from HuBERT-Large [16] [^5] We use checkpoint https://huggingface.co/facebook/hubert-large-ls960-ft. on the audio samples and their ground truth transcripts. We also report MOS-LQO (Mean Opinion Score – Listening Quality Objective) score to measure the reconstruction quality using ViSQOL [15]. Finally, we use human evaluation to measure the naturalness of the reconstructed speech using a MOS score with 95% confidence interval. We instruct the human evaluators to rate the naturalness of the reconstructed speech on a scale of 1 to 5, where 1 is the worst and 5 is the best. The results are shown in Table 3.
Table 3: Speech reconstruction results on 300 random samples from LibriSpeech test-clean set.
@@ 3.3.1 Speech Reconstruction
From Table 3, we can see that semantic tokenization (HuBERT-KM) is not good for speech reconstruction. Meanwhile, acoustic tokenizers that are optimized to reconstruct the signal directly (EnCodec and SpeechTokenizer) do well.
We apply different vocoders to reconstruct the speech signal from log mel-filterbanks, and find that the WER of the reconstructed speech signal is comparable to the acoustic tokenization methods with a fraction of the parameters. Also, log mel-filterbanks achieve a better MOS-LQO score, which indicates that the reconstructed audio is more similar to the original audio. By comparing Mel and dMel, we can see that discretization has little impact on WER and MOS-LQO scores. We also find that the exact vocoder matters much less than the frame rate of tokenization: the WER goes from 2.08 to 2.13 when switching from HifiGAN to ParallelWaveGAN, but it falls from 2.13 to 2.36 when the frame rate is changed from 80Hz to 40Hz. However, even a 1M parameter vocoder operating at a 40Hz frame rate is comparable to the much larger SpeechTokenizer on WER and MOS-LQO metrics.
@@
3.3.2 LM-Style Text-to-Speech
@@
3.3.3 LM-Style Speech-to-Text
@@
Training an LM-style speech-to-text (ASR) model can test if the speech tokens can preserve the semantic information in the speech signal and support the speech content-based task. Table 7 shows results of our model dubbed RichASR, trained with different tokenizations including dMel for the ASR task. Our LM-style model with dMel speech tokenization achieves 4.2% WER on the test-clean and 10.4% WER on the test-other sets outperforming both HuBERT-KM and SpeechTokenizer. We also observe that our model with HuBERT-KM [22] outperforms the SpeechTokenizer [46] for ASR, which is reasonable as semantic tokens are more suitable for the ASR task.
@@
3.3.4 Ablations
We first investigate the impact of the codebook sizes, shown in Table 9. The 16-bin configuration used in the paper demonstrates the best overall performance across tasks. While the 32-bin setup slightly outperforms on the ASR test-other set, it shows degraded performance in TTS. This trade-off likely stems from the increased speech vocabulary size, which may pose challenges for accurate prediction. The results may get better with increased data and model size. And 8-bin configuration looses too much information with discretization.
We then ablate the ASR results to understand why ASR LM-style model is behind the state-of-the-art ASR results on LibriSpeech. We take two existing transformer ASR baselines, Seq2Seq and CTC, that use 80 log mel-filterbanks and characters as targets. We then modify these baselines by using dMel instead (the discretization, embedding layer and linear layer) while keeping all other hyper-parameters the same (we adjust only the SpecAugment time masking max width accordingly to keep total masking in ms the same). Our results (Table 10) suggest: i) dMel brings only small degradation compared to Mel; ii) additional discrepancy is coming from different hop distance in featurization; iii) the main and significant performance degradation is coming from switching to LM-style model. The latter is in line with [24] and [10], though was not discussed in detail by any prior work. We hypothesise this gap is due to observed overfitting of the LM-style models.
Table 9: ASR and TTS results (WER, %) with dMel speech tokenizer and different number of bins (codebook size) for discretization in dMel. All models are trained on LibriSpeech 960h.
N-Bins ASR test-clean↓ ASR test-other↓ TTS WER↓ (%)
8 6.6 16.5 7.3
16 4.4 10.7 4.8
32 4.7 10.2 5.7
Model Features dev-clean↓ dev-other↓
[14] (RNN-T – Conformer) Mel-10ms 1.9 4.1
[20] (CTC – Squeezeformer) Mel-10ms 2.3 5.8
Seq2Seq [12] Mel-10ms 2.4 5.4
dMel-10ms 2.5 5.9
Mel-25ms 2.8 6.5
dMel-25ms 2.7 6.2
CTC [13] Mel-10ms 2.1 5.4
dMel-10ms 2.1 5.6
Mel-25ms 2.1 5.4
dMel-25ms 2.3 6.1
LM-style dMel-25ms 3.4 9.5
3.3.5 Unlocking Joint Speech-Text Modeling

### v3

@@ Abstract
Large language models have revolutionized natural language processing by leveraging self-supervised pretraining on vast textual data. Inspired by this success, researchers have investigated various compression-based speech tokenization methods to discretize continuous speech signals, enabling the application of language modeling techniques to discrete tokens. However, audio compressor introduces additional complexity and computational cost, and often fail on out-of-domain audio signals. In this work, we introduce a novel speech representation (dMel) that discretizes mel-filterbank channels into intensity bins, creating a simpler yet more effective representation compared to existing speech tokenization methods. Our approach demonstrates superior performance in preserving audio content, robustness to out-of-domain data, and offers a training-free, natural, and streamable representation. To address the high-dimensional nature of log-mel spectrograms, we propose an efficient parallel encoding and decoding method for high-dimensional tokens using an LM-style transformer architecture. This innovation enables us to develop RichTTS and RichASR—two models sharing the same architecture while achieving comparable or better results than specialized existing methods. Our results demonstrate the effectiveness of dMel in achieving high performance on both speech synthesis and recognition tasks within a unified framework, paving the way for efficient and effective joint modeling of speech and text. The code is available at https://github.com/apple/dmel and demos are available at https://apple.github.io/dmel-demo/.
@@ 1 Introduction
Current speech tokenization approaches can be broadly categorized into two types: semantic (content) tokens and acoustic tokens [^1] We use a word ‘semantic’ with the meaning of ‘content’ to keep prior work notation [8].. Semantic tokens, extracted from self-supervised (SSL) pretrained speech models [2, 18], are obtained by first encoding the speech signal into representations and then clustering them into discrete tokens with $k$-means method. However, such SSL pretrained models are not useful for high fidelity speech synthesis as speaker identity and other details of raw speech are lost in training [8]. Conversely, acoustic tokens can be obtained from audio compression models that are trained to compress the speech signal into codebook indices with residual vector quantization (RVQ) and reconstruction objectives [48, 12]. These tokens prioritize acoustic reconstruction but lose semantic information which can lead to poorer results in generating audio [44].
@@
We find that the word error rate (WER) of an ASR system run on the reconstructed waveforms is quite similar to the WER of the same system run on the ground-truth audio, showing that dMel captures the acoustic information needed to reconstruct good waveforms. Similarly, we find that the WER of ASR models trained on Mel and dMel are similar, indicating that dMel effectively preserves semantic content. This demonstrates that discretizing Mel has limited impact on information content while providing the benefits of discrete tokens.
@@
By operating on the log mel-filterbanks and preserving the frequency and intensity information (with some loss of resolution from discretization), dMel inherently preserves both semantic and acoustic information in a unified representation, without the need for separate tokenization or additional pretraining of a tokenization model. The key advantages of dMel include:
  • Interpretable and complete: Preserves both semantic and acoustic information in an interpretable physics-based representation, with minimal information loss from discretization.
  • Model-free and versatile: Directly compatible with any mel-filterbank vocoder for waveform reconstruction, unlike other methods where representations are tightly coupled to specific encoder-decoder architectures.
  • Parallel processing: Frequency channels can be modeled independently without complex hierarchical dependencies, enabling efficient processing with decoder-only transformer architectures.
While data-driven tokenization methods can be improved with more diverse training data, they inherently suffer from information loss due to their neural compression—discarding information that may be crucial in unseen conditions. In contrast, mel-spectrogram offers advantages as a physics-based signal representation:
  • Robust frequency preservation: Captures frequency components through principled transformation, aligning with human auditory perception’s sensitivity to magnitude spectrum.
  • Domain-agnostic performance: Demonstrates consistent behavior across acoustic conditions without requiring domain-specific training.
  • Proven reliability: Shows robust performance across decades of speech processing research in diverse conditions.
Given the high-dimensional nature of log-mel spectrograms, we propose an efficient parallel encoding and decoding method for these high-dimensional tokens using an LM-style transformer architecture. This enables us to develop RichTTS and RichASR—two models sharing the same architecture while achieving comparable or better results than specialized existing methods.
@@
Through comprehensive evaluations, we show that using dMel allows us to employ a single decoder-only model to achieve high performance on both automatic speech recognition ASR and TTS tasks. The ASR task validates that dMel preserves semantic information, while the TTS task demonstrates that dMel is effective for high-fidelity acoustic reconstruction of speech. We also compare dMel to other tokenization methods and find that dMel achieves the best WER for the ASR task, which indicates that semantic information is well preserved. Additionally, dMel achieves a lower WER score for the TTS task when using WhisperX [4] for automatic evaluation, and we find that models trained with dMel can generate long and natural speech samples.
@@
Figure 1: Prior works on speech tokenization use either heavy self-supervised pretrained encoders [2, 18] to extract semantic tokens (and train a separate decoder for it [24]) or learn compression encoder-decoder models with residual vector quantizations [48, 12] to obtain acoustic tokens. By contrast we eliminate the encoder and simply discretize mel-filterbanks (dMel) to encode audio, and use a simple mel-filterbank vocoder [47] to reconstruct speech signals.
@@ 2 Method
Comparison between Speech Tokenizers
Table 2: Comparison between different speech tokenizers: dMel (ours), HuBERT-KM and SpeechTokenizer. For dMel we use $N=80$ log mel-filterbanks (50ms window, 25ms hop distance), and $162^{K}=$ 16 values of the codebook $\mathbf{C}$. For HuBERT-KM, 200 is chosen accoring to [28].
dMel HuBERT-KM SpeechTokenizer
Codebook Size 16 200 1024
Code Dimension 80 1 8
Vocabulary Size 16 * 1 200 * 1 1024 * 8
Frame-rate 40Hz 50Hz 50Hz
Bit-rate 12.8kps 0.4kps 4kps
Training-free?
In Table2, we compare dMel with the baselines in terms of vocabulary size, bit-rate, and frame-rate. First, dMel has a much smaller vocabulary, as it is discretized mel-filterbanks energies, allowing all 80 channels to share the same vocabulary since they represent similar energy values. In contrast, neural compression encoders like SpeechTokenizer require separate embeddings for different channels. Also, dMel operates at a lower frame-rate while maintaining a higher bit-rate. The reduced frame-rate leads to shorter sequence lengths during both training and inference, which is particularly advantageous when using large models. While a higher bit-rate typically increases model complexity for compression-based tokenizers, this is not the case for dMel due to two key factors: i) dMel is encoder-free, without any compression encoder; ii) the complexity of the model introduced in Section 2.2 depends only on the vocabulary size and sequence length (frame-rate), not on the code dimensions. In compression-based methods, increasing the bit-rate requires either larger codebooks or additional residual dimensions, leading to increased tokenizer complexity. Moreover, these methods require more complex downstream models to handle the expanded representations. Given recent studies [13, 30] demonstrated that bit-rate does not strongly correlate with downstream model performance, we focus our comparative analysis on frame-rate rather than bit-rate when evaluating different speech tokens for downstream tasks. This approach challenges the conventional assumption that higher bit-rates necessarily yield better results.
2.2 Unified Speech-Text Transformer Decoder (RichTTS and RichASR)
Figure 2: (Left) For a time step $t$ encoded dMel from Figure 1 is inputted to the transformer decoder to produce final embeddings for each of the frequency channels in parallel. (Right) Unified Speech-Text Transformer Decoder with speech tokens as dMel.
@@ 2.2 Unified Speech-Text Transformer Decoder (RichTTS and RichASR)
Figure 3: Unified Speech-Text Transformer Decoder with speech tokens as dMel where we predict multiple, e.g. two, frames in parallel reducing the frame rate, e.g. by 2x: dMel tokens for every two frames are stacked together to form the input into the decoder and predicted in parallel afterwards.
@@
The transformer decoder is trained end-to-end on a combined dataset of speech and text pairs. For TTS training, the input sequence is constructed by concatenating the speaker embedding (The speaker embeddings are extracted from an independent dvector [43] model [^3] We use a pretrained model “Speaker Encoder” from the YourTTS [10] repository https://github.com/Edresson/YourTTS.), text tokens, and speech tokens. For ASR training, the input sequence is constructed by concatenating the speech tokens and text tokens. Both tasks are trained with causal masking, where the model is trained to predict the next token based on the previous tokens. The loss is calculated using the cross-entropy loss between the predicted tokens and the ground-truth tokens. Loss calculation is skipped on the speech tokens for ASR task and on the text tokens for TTS task. Note, that all frequency channels at time frame $t$ for dMel tokenizer are predicted independently and in parallel, see Figure 2 (left). For positions encoding, to capture the relative distances between tokens in the input sequence, we apply multiplicative relative positional embedding RoPE [41]. This allows the model to learn the positional relationships between speech tokens, text tokens, and speaker embeddings, enhancing its ability to generate coherent output sequences. For positional embeddings we do not distinguish between text, speech and speaker tokens and thus having global positions notation across all of them, see Figure 2.
Context Masking during Training
@@
k-Frame Encoding and Decoding
Given the fact that our model can encode and decode multiuple channels in parallel, for mel-spectrogram, we are insterested in whether the model can do k-frame encoding and decoding in parallel too. As shown in Figure 3, we simply concatenate multiple frames together and using the same technique to encode and decode it as described in previous paragraphs. Our results show that this model is also working well on multi-frame generation in parall. More results can be found in Section 3.
@@ 3 Experiments
3.1 Speech Reconstruction on Clean Speech
We first conduct speech reconstruction experiments with various speech tokenizers on clean speech. Following [52], we randomly sample 300 speech utterances and their ground truth transcriptions from the LibriSpeech test-clean dataset. We use the speech2unit and unit2speech modules to convert the speech signal to speech tokens and then reconstruct the speech signal from the speech tokens. We compute the WER between the ASR outputs from HuBERT-Large [18] [^4] We use checkpoint https://huggingface.co/facebook/hubert-large-ls960-ft. on the audio samples and their ground truth transcripts. We also report MOS-LQO (Mean Opinion Score – Listening Quality Objective) score to measure the reconstruction quality using ViSQOL [17]. Finally, we use human evaluation to measure the naturalness of the reconstructed speech using a MOS score with 95% confidence interval. We instruct the human evaluators to rate the naturalness of the reconstructed speech on a scale of 1 to 5, where 1 is the worst and 5 is the best. The results are shown in Table 3.
Table 3: Speech reconstruction results on 300 random samples from LibriSpeech test-clean set. WER (%) is evaluated with Hubert Large.
@@ 3.1 Speech Reconstruction on Clean Speech
Figure 4: Speech reconstruction results on 300 random samples from LibriSpeech test-clean set when noise is added: either background music from [7] dataset or speech noise from test-other. WER (%) is evaluated with WhisperX ASR (“base.en”). Audio examples are in our demo page.
From Table 3, we can see that semantic tokenization (HuBERT-KM) is poor for speech reconstruction. Meanwhile, acoustic tokenizers that are optimized to reconstruct the signal directly (EnCodec and SpeechTokenizer) do well.
We apply different vocoders to reconstruct the speech signal from log mel-filterbanks, and find that the WER of the reconstructed speech signal is comparable to the acoustic tokenization methods with a fraction of the parameters. Also, log mel-filterbanks achieve a better MOS-LQO score, which indicates that the reconstructed audio is more similar to the original audio. By comparing Mel and dMel, we can see that discretization has little impact on WER and MOS-LQO scores. We also find that the exact vocoder matters much less than the frame rate of tokenization: the WER goes from 2.08 to 2.13 when switching from HifiGAN to ParallelWaveGAN (PWG), but it falls from 2.13 to 2.36 when the frame rate is changed from 80Hz to 40Hz. However, even a 1M parameter vocoder operating at a 40Hz frame rate is comparable to the much larger SpeechTokenizer on WER and MOS-LQO metrics.
3.2 Speech Reconstruction on Noisy Speech
In addition to the in-domain clean human speech, we also measure speech tokenizer’s out-of-domain ability which is neglected in prior work. We compare different speech tokenizations for speech reconstruction under noisy conditions. We syntetically add two kinds of noise signal to the clean audio: i) add music background to the original human speech; ii) add another human speech in low volume as a background to the original human speech. Results are shown in Figure 4: both HuBERT-KM and SpeechTokenizer fail in out-of-domain setting while EnCodec, Mel and dMel show robustness for noisy speech reconstruction. This supports our motivation to explore dMel, training-free and determenistic tokenization, which is able to handle various acoustic conditions.
@@ 3.2 Speech Reconstruction on Noisy Speech
3.3 LM-Style Text-to-Speech
@@
3.3.1 Configurations
We use several open-sourced datasets with paired speech and text transcription to conduct experiments: i) LibriSpeech [31] dataset (CC BY 4.0) consists of English speech recordings (960h, 16kHz) from various speakers ($\sim$2k) and conditions; ii) LibriTTS [49] (CC BY 4.0) dataset (500h) derived from LibriSpeech improves on it with the proper sentence split, text normalization and keeping samples 24kHz; iii) VCTK [46] contains 44h of English speech (108 speakers); iv) LJSpeech [19] (public domain in US) is a single speaker English audio recordings of 22kHz with read speech from LibriVox [^5] https://librivox.org/pages/public-domain/.. While LibriSpeech is used to train ASR and TTS models, LibriTTS, VCTK and LJSpeech are only used to train the TTS.
@@ 3.3.2 Results
3.4 LM-Style Speech-to-Text
@@
Training an LM-style speech-to-text (ASR) model can test if the speech tokens can preserve the semantic information in the speech signal and support the speech content-based task. We use the same experiments configuration in Section3.3.1 to train ASR models. Table 7 shows results of our model dubbed RichASR, trained with different tokenizations including dMel for the ASR task. Our LM-style model with dMel speech tokenization achieves 4.2% WER on the test-clean and 10.4% WER on the test-other sets outperforming both HuBERT-KM and SpeechTokenizer. We also observe that our model with HuBERT-KM [24] outperforms the SpeechTokenizer [52] for ASR, which is reasonable as semantic tokens are more suitable for the ASR task.
@@
3.5 Ablations
We first investigate the impact of the codebook sizes, shown in Figure 6. The 16-bin configuration used in the paper demonstrates the best overall performance across tasks. While the 32-bin setup slightly outperforms on the ASR test-other set, it shows degraded performance in TTS. This trade-off likely stems from the increased speech vocabulary size, which may pose challenges for accurate prediction. The results may get better with increased data and model size. And 8-bin configuration looses too much information with discretization.
We then ablate the ASR results to understand why ASR LM-style model is behind the state-of-the-art on LibriSpeech. We take two existing transformer ASR baselines, Seq2Seq and CTC, that use 80 log mel-filterbanks and characters as targets. We then modify these baselines by using dMel instead (the discretization, embedding layer and linear layer) while keeping all other hyper-parameters the same (we adjust only the SpecAugment time masking max width accordingly to keep total masking in ms the same). Our results (Appendix Table 12) suggest: i) dMel brings only small degradation compared to Mel; ii) additional discrepancy is coming from different hop distance in featurization; iii) the main and significant performance degradation is coming from switching to LM-style model. The latter is in line with [28] and [11], though was not discussed in detail by any prior work. We hypothesise this gap is due to observed overfitting of the LM-style models.
Finally, we conduct ablation on k-frame parallel encoding and decoding introduced in Section 2.2. The results are shown in Figure 6. As we can see from this figure, $k<=4$ yeild similar results to single-frame model, while improves both the training and inference efficiency significantly. In contrast, for SpeechTokenizer, even $K=1$ is worse than dMel with $k=6$. This is because the residual nature of VQ tokens in SpeechTokenizer where each channel is a residual of the previous channel, which is not suitable for multi-channel parallel decoding.
Figure 5: ASR and TTS results (WER, %) with dMel speech tokenizer and different number of bins (codebook size) for discretization in dMel. All models are trained on LibriSpeech 960h.
Figure 6: Ablation results on k-frame encoding and decoding.
3.6 Unlocking Joint Speech-Text Modeling